# RegGS: Unposed Sparse Views Gaussian Splatting with 3DGS Registration

### v1

@@ 1 Introduction
To address this issue, we propose a novel 3D Gaussian reconstruction framework: RegGS, which performs unposed sparse-view reconstruction by registering feed-forward Gaussian incrementally. Specifically, we introduce the optimal transport-based Mixture 2-Wasserstein $(MW2)$ distance between Gaussian mixture models (GMM) to align generalized Gaussian manifolds. Through a differentiable multi-modal joint registration pipeline, we solve for scene alignment in the $\mathrm{Sim(3)}$ space.
@@ 2.1 NeRF-based Pose-Free Reconstruction
Several approaches have been proposed to extend NeRF to handle unposed input images. [8, 12] integrates camera pose estimation with NeRF rendering, leveraging a recurrent GRU module for pose and depth estimation. Similarly, [36] employs a weighted Procrustes analysis and an optical flow network to establish correspondences for pose estimation. More recently, CoPoNeRF [22] introduced a unified framework that integrates correspondence matching, pose estimation, and NeRF rendering, allowing for end-to-end training and improved performance in challenging scenarios with extreme viewpoint changes. Additionally, methods like Nope-NeRF [4] leverage depth information to constrain the optimization process.
@@ 2.3 Feedforward-based Pose-Free 3DGS Reconstruction
Consequently, we explored a method based on 3D Gaussian registration to achieve incremental unposed sparse-view reconstruction. This approach not only leverages the excellent scene priors of feed-forward models but also enables high-quality reconstruction in broader sparse view scenarios, which is of practical importance.
@@ 4.1 Experiment Setup
For unposed sparse views reconstruction task, the number of views we reconstructed are 2, 8, 16, and 32. To simulate sparse input, both training and testing views are equidistantly sampled from the videos. For 2-view scenarios, we sample every 40 frames for videos with significant motion and every 60 frames for scenes with less motion. For scenarios with 8, 16, and 32 views, training views are equidistantly sampled throughout the entire video. The test set includes all frames not used for training.
@@ 4.2 Experimental results and Analysis
Novel View Synthesis: As shown in Tab. 1, and Tab. 2, and Fig. 3 and Fig. 4, our method significantly outperforms other unposed reconstruction methods in terms of PSNR and SSIM. NoPe-NeRF [4] fails to converge; VideoLifter [14] produces distorted renderings under sparse views; and CF-3DGS [20] suffers from artifacts due to inadequate detail capture. On the LPIPS, we generally lead, though we occasionally fall short in some cases, due to noise introduced by global refinement when improving PSNR.
Pose Estimation: Our method can also be applied to pose estimation. As shown in Tab. 3 and Fig. 6. We conducted experiments on the RE10K [46] and ACID [26] datasets under 8, 16, and 32-view input conditions. The poses estimated by the baseline methods were aligned with the ground truth (GT) poses for comparison. Table 3 presents the performance of our method. Compared to other unposed methods, our method demonstrates a more pronounced performance gap, especially in sparse view conditions.
@@ 5 Conclusion
This paper presents RegGS, an incremental 3D Gaussian reconstruction framework for unposed sparse-view settings. We constructed a GMM alignment metric in $\mathrm{Sim}(3)$ space based on the optimal transport $\text{MW}_{2}$ distance, and efficiently computed the $\text{MW}_{2}$ distance using the entropy-regularized Sinkhorn algorithm, thereby circumventing the infinite-dimensional optimization problem. By jointly optimizing $\text{MW}_{2}$, photometric, and depth-consistency losses, RegGS achieves progressive coarse-to-fine registration of both camera poses and scene structure. Experiments on RE10K and ACID demonstrate superior pose estimation and novel view synthesis compared to prior methods, highlighting RegGS’s potential for real-world applications.
@@ 6.3 Construction of the Subgradient Set
7 Sinkhorn Algorithm Complexity Calculation
@@ 8 Additional Experimental Results
We evaluate runtime and memory across sampling ratios from 2x to 64x using 200-frame sequences. As shown in Table 7, RegGS maintains controlled memory usage across all input settings. In contrast, NoPoSplat and Splatt3R are limited to two-view inputs, while DUSt3R and MASt3R exhibit exponential growth in Gaussian count, frequently resulting in out-of-memory failures. This demonstrates the scalability of RegGS under sparse-view conditions. At 64x, as view coverage becomes denser, the reconstruction bottleneck shifts from view sparsity to the capacity of the 3DGS representation. RegGS achieves comparable reconstruction quality to MASt3R with significantly lower memory consumption. Further optimization of $\text{MW}_{2}$ computation remains a direction for future work.

### v2

@@ 1 Introduction
To address this issue, we propose a novel 3D Gaussian reconstruction framework: RegGS, which performs unposed sparse view reconstruction by registering feed-forward Gaussian incrementally. Specifically, we introduce the optimal transport-based Mixture 2-Wasserstein $(MW2)$ distance between Gaussian mixture models (GMM) to align generalized Gaussian manifolds. Through a differentiable multi-modal joint registration pipeline, we solve for scene alignment in the $\mathrm{Sim(3)}$ space.
@@ 2.1 NeRF-based Pose-Free Reconstruction
Several approaches have been proposed to extend NeRF to handle unposed input images. Among them, [8, 12] integrate camera pose estimation with NeRF rendering, leveraging a recurrent GRU module for pose and depth estimation. Similarly, [36] employs a weighted Procrustes analysis and an optical flow network to establish correspondences for pose estimation. More recently, CoPoNeRF [22] introduced a unified framework that integrates correspondence matching, pose estimation, and NeRF rendering, allowing for end-to-end training and improved performance in challenging scenarios with extreme viewpoint changes. Additionally, methods like Nope-NeRF [4] leverage depth information to constrain the optimization process.
@@ 2.3 Feedforward-based Pose-Free 3DGS Reconstruction
Consequently, we explored a method based on 3D Gaussian registration to achieve incremental unposed sparse view reconstruction. This approach not only leverages the excellent scene priors of feed-forward models but also enables high-quality reconstruction in broader sparse view scenarios, which is of practical importance.
@@ 4.1 Experiment Setup
For the unposed sparse views reconstruction task, the number of views we reconstructed are 2, 8, 16, and 32. To simulate sparse input, both training and testing views are equidistantly sampled from the videos. For 2-view scenarios, we sample every 40 frames for videos with significant motion and every 60 frames for scenes with less motion. For scenarios with 8, 16, and 32 views, training views are equidistantly sampled throughout the entire video. The test set includes all frames not used for training.
@@ 4.2 Experimental Results and Analysis
Novel View Synthesis: As shown in Tab. 1, Tab. 2, Fig. 3 and Fig. 4, our method significantly outperforms other unposed reconstruction methods in terms of PSNR and SSIM. NoPe-NeRF [4] fails to converge; VideoLifter [14] produces distorted renderings under sparse views; and CF-3DGS [20] suffers from artifacts due to inadequate detail capture. For LPIPS, we generally lead, though we occasionally fall short in some cases, due to noise introduced by global refinement when improving PSNR.
Pose Estimation: Our method can also be applied to pose estimation. As shown in Tab. 3 and Fig. 6. We conduct experiments on the RE10K [46] and ACID [26] datasets under 8, 16, and 32-view input conditions. The poses estimated by the baseline methods were aligned with the ground truth (GT) poses for comparison. Table 3 presents the performance of our method. Compared to other unposed methods, our method demonstrates a more pronounced performance gap, especially in sparse view conditions.
@@ 5 Conclusion
This paper presents RegGS, an incremental 3D Gaussian reconstruction framework for unposed sparse view settings. We constructed a GMM alignment metric in $\mathrm{Sim}(3)$ space based on the optimal transport $\text{MW}_{2}$ distance, and efficiently computed the $\text{MW}_{2}$ distance using the entropy-regularized Sinkhorn algorithm, thereby circumventing the infinite-dimensional optimization problem. By jointly optimizing $\text{MW}_{2}$, photometric, and depth-consistency losses, RegGS achieves progressive coarse-to-fine registration of both camera poses and scene structure. Experiments on RE10K and ACID demonstrate superior pose estimation and novel view synthesis compared to prior methods, highlighting RegGS’s potential for real-world applications.
@@ 6.3 Construction of the Subgradient Set
7 Sinkhorn Algorithm Complexity
@@ 8 Additional Experimental Results
We evaluate runtime and memory across sampling ratios from 2x to 64x using 200-frame sequences. As shown in Tab. 7, RegGS maintains controlled memory usage across all input settings. In contrast, NoPoSplat and Splatt3R are limited to two-view inputs, while DUSt3R and MASt3R exhibit exponential growth in Gaussian count, frequently resulting in out-of-memory failures. This demonstrates the scalability of RegGS under sparse view conditions. At 64x, as view coverage becomes denser, the reconstruction bottleneck shifts from view sparsity to the capacity of the 3DGS representation. RegGS achieves comparable reconstruction quality to MASt3R with significantly lower memory consumption. Further optimization of $\text{MW}_{2}$ computation remains a direction for future work.